OpenStudy (anonymous):

4sinxcosx+2sinx-2cosx-1=0

OpenStudy (anonymous):

are u wanting to solve

OpenStudy (anonymous):

@cxxonicole98 what shoul i hv to do

OpenStudy (anonymous):

what are you asking? @rizzu

OpenStudy (anonymous):

@cxxonicole98 i am asking about que........i thnk we hv to find the value of x in this que

OpenStudy (welshfella):

4sinxcosx+2sinx-2cosx-1=0 2 sin x(2 cos x + 1) -(2 cos x + 1) = 0 (2 sin x - 1)(2 cos x = 1) = 0

OpenStudy (welshfella):

* that's 2 cos x + 1 in the second parentheses

OpenStudy (welshfella):

now you have 2 sin x - 1 = 0 and 2 cos x + 1 = 0 solve these 2 equations and you are done

OpenStudy (welshfella):

find what cos x and sin x are then you can find the values of x I suggest you find all values between 0 and 360 degrees